# Sensitivity of CTAO to axion-like particles from blazars: a machine learning approach

### v1

@@ 4.2 Method
In the following we introduce our technique to derive upper limits on ALPs by using ML algorithms to classify binned gamma-ray datasets. After being trained with a large amount of simulated spectra, both with and without ALP effects, the classifier is supposed to learn what the most important energy bins are to distinguish between the two classes, i.e., where the spectral oscillations and/or hardening are most relevant. This classification task crucially depends on the source under consideration, as well as on the ALP parameters $(m_{a},g_{a\gamma})$. Our approach was loosely inspired by that described in ref. [day2020].
@@
These counts are then normalised to the $[0,1]$ interval, so that the only relevant characteristic for classifying the spectra is their shape, given by the intrinsic differences between adjacent bins rather than the overall number of counts. In order to allow for some variability in the intrinsic spectrum, avoiding to overfit the classifiers to a specific model, the training datasets are simulated by uniformly sampling values of each spectral parameter $p$ in the range $p-\Delta p,p+\Delta p]$, according to the values reported in Table 3. We have checked that the classifier performance only improves slightly by further increasing the number of training datasets.
@@
    n_estimators=100, max_depth=3, reg_lambda=500., reg_alpha=10., eta=0.3
While this choice is arbitrary to a certain extent, we computed the values of several metrics (accuracy, precision, recall, F1 score) over the grid to check that the performance is reasonably good in the same regions of the parameter space where we expect ALP signatures to be stronger. Indeed, a decrease in performance is observed at large $m_{a}$ and small $g_{a\gamma}$ values, in the lower right corner of the grid, where the spectral distortions due to ALPs are weaker and less distinguishable from the ALP-less case. As an example, the results obtained for the baseline state of Mrk 501 are shown in Figure 4; see also Appendix B. These metrics were computed with Scikit-learn, using a 80-20 train-test split. For their definitions, see e.g. ref. [raschka2017].
Figure 4: XGB metrics for the baseline state of Mrk 501 over the ALP parameter space.
@@
Figure 5: Example distributions of the $\Pi$ statistic defined in eq. (4.1), comparing good and suboptimal performance of a classifier algorithm in different regions of the ALP parameter space.
We then search a $\Pi$ distribution for each $(m_{a},g_{a\gamma})$ pair, evaluating the $\Pi$ statistic in eq. (4.1) for 2000 simulated ALP-like datasets. Since the value of $\Pi$ is always between 0 and 1, we fit the resulting distribution to a Beta PDF
$f(x,\alpha,\beta)=\frac{\Gamma(\alpha+\beta)x^{\alpha-1}(1-x)^{\beta-1}}{\Gamma(\alpha)\Gamma(\beta)}\,$, (4.2)
with $0\leq x\leq 1$ and $\alpha,\beta>0$. Using this distribution, the confidence level for excluding the ALP hypothesis and the $z$-score are computed using eqs. (3.8) and (3.9), respectively. The quality of the fit is generally good for medium and low values of $g_{a\gamma}$, as can be seen from Figure 5, but it worsens for higher ones. We have checked in Appendix C that this does not significantly impact the main conclusions of this work.
@@ 5 Discussion of results
Our main results are presented in Figure 6, where a color map is used to represent the exclusion significance obtained with the ML classifier method in units of Gaussian standard deviations over the ALP parameter space. We note that, for both methods, the exclusion significance may be infinite for those points where the confidence level computed by eq. (3.8) is equal to 1. For plotting, these values are clipped to the highest finite significance; see also Appendix C.
Figure 6: ALP exclusion significance computed with the ML-based method, with the $2\sigma$ contours highlighted (solid lines). The contours obtained at the same confidence level with the LRT method are shown as dashed lines for comparison. A linear interpolation was applied between the points in the parameter space for which the exclusion confidence level was computed, which are indicated as black dots.
@@
The $2\sigma$ sensitivity obtained with the method of ML classifiers, shown as solid lines in Figure 6, is comparable with that resulting from the LRT (dashed lines). For all sources, the overall shapes and extensions of the projected exclusion regions are in agreement between the two methods, with occasional low-significance “holes” or high-significance “islands” in both sets of contours. Such features are common to other analyses of this kind [ajello2016, abdalla2021] and do not significantly affect our main results.
@@
The excluded region of the parameter space in the case of Mrk 501 is almost the same for both the baseline and flaring states. An explanation for this is the fact that the spectral shapes considered for the two states are fairly similar (see Fig. 1), and that the increased brightness of the flaring state is compensated by the longer exposure used for the baseline state. Given this, we can assume that the photon counts in the considered energy ranges are above the CTAO sensitivity for both states and are equally affected by ALPs, thus leading to similar exclusions in the parameter space. Conversely, PKS 2155-304 exhibits a spectral curvature and a stronger EBL absorption effect that significantly reduce the flux at higher energies. In this case, a flaring state can indeed bring the flux of higher-energy photons above the CTAO sensitivity, thus improving the constraining power of the observations. These qualitative considerations are confirmed by computing the Li-Ma significance $S$ of the signal [li1983]. While for Mrk 501 both observations yield $S\sim 300$, in the case of PKS 2155-304 a 5-hour flaring observation results in a significance $S\sim 800$, compared to $S\sim 200$ for a 50-hour observation in the baseline state.
@@
These projections, representing the mean expected CTAO sensitivity to ALPs from each considered source, are obtained by averaging the results from many different simulated observations. A different approach for estimating such a mean sensitivity, employed in various studies on ALPs [meyer2014a, meyer2014b, abdalla2021] and WIMPs [acharyya2021, acharyya2023, abe2024b], would be to use so-called “Asimov datasets” [cowan2011] instead. These datasets are generated by simply setting the numbers of counts in the ON and OFF regions equal to their expected values, in order to eliminate Poissonian fluctuations and directly provide a median estimate of the sensitivity. We have checked explicitly that the contours obtained from a single Asimov dataset using the LRT method are very close to those resulting from averaging over many “noisy” datasets. On the other hand, we found that applying the ML classifiers to Asimov datasets causes a significant overestimation of the exclusion regions, as the total absence of Poissonian noise in photon counts almost always produces low $p_{\rm ALP}$ values even for extremely small $g_{a\gamma}$. In a realistic analysis, these points of the parameter space would not be excluded, as the inevitable photon noise could easily mimic ALP effects at very low couplings. We thus conclude that Asimov datasets cannot be reliably used in estimating ALP sensitivities with ML classifiers.
@@
As a final remark, we note that these results may be affected by several systematic effects, such as the choice of energy binning, magnetic field parameters and EBL model. We discuss these effects in Appendix D. We note, in particular, that recent measurements of the cosmic infrared background by the CIBER experiment might extend the allowed parameter space for ALPs to a wider region than what is expected from the commonly adopted EBL models [kohri2017].
@@ 6 Conclusions
In this work, we have estimated the capability of CTAO to place upper limits on the ALP parameter space by simulating observations of the blazars Mrk 501 and PKS 2155-304 with the Northern and Southern Arrays of the Observatory. The source selection and observation criteria were motivated by the AGN Key Science Project described in ref. [ctascience2018]. Mock datasets for each source were prepared considering both long-term observations of baseline states, with spectra taken from the Fermi-LAT 4FGL-DR4 source catalog [4fgl-dr4], and short flares based on previous TeV observations [albert2007, aharonian2009].
We have proposed a new computational approach to constrain the ALP parameter space with binned astrophysical gamma-ray data, described in Section 4, based on the usage of machine learning classifiers. In Figure 6 we show the mean exclusion significance that can be obtained by applying this technique to our simulated datasets. The sensitivity at a confidence level of $2\sigma$ is comparable to that resulting from standard likelihood-ratio test, suggesting that our approach is a valid alternative to more traditional analysis methods. We note, however, that our results are based on a somewhat arbitrary choice of the classifiers’ hyperparameters. A more systematic tuning may be required to obtain more accurate predictions, and to generalize or even enhance the performance of this technique.
Figure 7: Comparison of the CTAO $2\sigma$ exclusion regions on the ALP parameter space obtained in this work with reference constraints from the literature (green). The projected sensitivities obtained from 50-hour observations of Mrk 501 (red) and PKS 2155-304 (blue) in their baseline states, as well as that from a 5-hour observation of a PKS 2155-304 flare (yellow) are shown. We do not explicitly include the results obtained from the Mrk 501 flare as they largely overlap with those from the baseline observation.
@@
In Figure 7 we compare the projections derived in this work with previously published constraints on the ALP parameter space [^11] The reference limits are mostly taken from the online repository at ref. [AxionLimits].. Most of these are computed at the $2\sigma$ confidence level [dessert2022, noordhuis2023, manzari2024, ajello2016, altenmuller2024, abdalla2021, abramowski2013], except for the limits by MAGIC which are reported at the $3\sigma$ level [abe2024]. In particular, ref. [abdalla2021] represents the only previous estimate of ALP sensitivity officially published by the CTAO Consortium, derived from simulated observations of the AGN NGC 1275 inside the Perseus galaxy cluster. With these targets, we expect that CTAO will be able to place limits on a much wider portion of previously unconstrained ALP parameter space in the $0.1\mbox{--}100\,\rm{neV}$ mass range. We finally remark that, while already competitive, the limits derived from a Southern source like PKS 2155-304 could be further improved with the foreseen addition of 2 LSTs to the Southern Array of CTAO, whose contribution would increase the sensitivity to low-energy spectral oscillations.
@@ Acknowledgments
This work was conducted in the context of the CTAO DMEP working group. CTAO gratefully acknowledges financial support from the agencies and organizations listed at
https://www.ctao.org/for-scientists/library/acknowledgments/.
@@
We would like to thank the computing centers that provided resources for the IRF generation, listed at https://zenodo.org/records/5499840.
Source Model $\phi_{0}$ $E_{0}$ $\alpha$ $\beta$ $E_{\rm cut}$ Ref. Baseline $10^{-12}\,\rm{MeV^{-1}\,cm^{-2}\,s^{-1}}$] [MeV] Mrk 501 LP $3.76\pm 0.05$ 1535.09 $1.75\pm 0.01$ $0.019\pm 0.004$ - [4fgl-dr4] PKS 2155-304 LP $2.59\pm 0.13 11$60.97 $1.77\pm 0.01$ $0.042\pm 0.004$ - [4fgl-dr4] Flaring $10^{-10}\,\rm{TeV^{-1}\,cm^{-2}\,s^{-1}}$] [TeV] [TeV] Mrk 501 LP $18.6\pm 0.06$ 0.3 $1.73\pm 0.06$ $0.13\pm 0.03$ - [albert2007] PKS 2155-304 ECPL $4.51\pm 0.36$ 1.0 $2.16\pm 0.05$ - $1.74\pm 0.24$ [aharonian2009]
Table 3: Parameters of the spectral models used for the sources considered in this work. We note that the $\beta$ parameter for the Mrk 501 flaring state is scaled by a factor $\ln(10)$ with respect to ref. [albert2007] after the conversion from base-10 logarithm to natural logarithm.
Appendix A Details on spectral models
@@
The parameters of each spectral model considered in this work are reported in Table 3, where LP and ECPL refer to the log-parabola and power law with exponential cutoff models defined in eqs. (2.1) and (2.2). These models were implemented in Gammapy using the LogParabolaSpectralModel and ExpCutoffPowerLawSpectralModel classes, respectively. As discussed in the main text, the baseline spectra were taken from the Fermi-LAT 4FGL-DR4 catalog [4fgl-dr4], while the flaring spectra were selected from previously published IACT observations. In particular, the EBL-corrected parameters of the Mrk 501 flaring spectrum are taken from Table 7 and Figure 19 of ref. [albert2007], referring to the flare of 30 June 2005. The PKS 2155-304 parameters are quoted from Table 3 in ref. [aharonian2009] and correspond to the data taken with the lowest energy threshold ($\rm 200\,GeV$, labelled as T200), corrected for EBL absorption.
@@
Figure 8: Accuracy of the XGBoost classifier grids defined over the ALP parameter space with the hyperparameters defined in Section 4, with the $90\%$ contours highlighted.
Appendix B Classifier accuracy over the ALP parameter space

### v2

@@ 4.2 Method
In general, the effects of ALP-photon conversions on gamma-ray spectra are highly nonlinear and not easily parametrized. The energy and amplitude at which spectral features appear depend in a nontrivial way from the morphology and intensity of the magnetic fields encountered by the ALP-photon beam along the line of sight from Earth. It is therefore interesting to see how non-parametric statistical methods can address this task, such as those provided by ML classification algorithms. A first step in this direction is to verify that a ML-based method can reproduce the results obtained by the classical likelihood-ratio test (LRT) discussed above.
After being trained with a large amount of simulated spectra, both with and without ALP effects, a classifier is supposed to learn what the most important energy bins are to distinguish between the two classes, i.e., where the spectral oscillations and/or hardening are most relevant. This classification task crucially depends on the source under consideration, as well as on the ALP parameters $(m_{a},g_{a\gamma})$. Our approach was loosely inspired by that described in ref. Day and Krippendorf (2020).
@@
These counts are then normalized to the $[0,1]$ interval, so that the only relevant characteristic for classifying the spectra is their shape, given by the intrinsic differences between adjacent bins rather than the overall number of counts. In order to allow for some variability in the intrinsic spectrum, avoiding to overfit the classifiers to a specific model, the training datasets are simulated by uniformly sampling values of each spectral parameter $p$ in the range $p-\Delta p,p+\Delta p]$, according to the values reported in Table 3. We have checked that the classifier performance only improves slightly by further increasing the number of training datasets.
@@
    n_estimators=100, max_depth=3,
    reg_lambda=500., reg_alpha=10.,
    eta=0.3
While this choice is arbitrary to a certain extent, we computed the values of several metrics (accuracy, precision, recall, F1 score) over the grid to check that the performance is reasonably good in the same regions of the parameter space where we expect ALP signatures to be stronger. Indeed, a decrease in performance is observed at large $m_{a}$ and small $g_{a\gamma}$ values, in the lower right corner of the grid, where the spectral distortions due to ALPs are weaker and less distinguishable from the ALP-less case. As an example, the results obtained for the baseline state of Mrk 501 are shown in Figure 4; see also Appendix C. These metrics were computed with Scikit-learn, using a 80-20 train-test split. For their definitions, see e.g. ref. Raschka and Mirjalili (2017).
Figure 4: XGB metrics for the baseline state of Mrk 501 over the ALP parameter space
@@
We then search a $\Pi$ distribution for each $(m_{a},g_{a\gamma})$ pair, evaluating the $\Pi$ statistic in eq. (19) for 2 000 simulated ALP-like datasets. Since the value of $\Pi$ is always between 0 and 1, we fit the resulting distribution to a Beta PDF
$f(x,\alpha,\beta)=\frac{\Gamma(\alpha+\beta)x^{\alpha-1}(1-x)^{\beta-1}}{\Gamma(\alpha)\Gamma(\beta)}\,$, (20)
with $0\leq x\leq 1$ and $\alpha,\beta>0$. Using this distribution, the confidence level for excluding the ALP hypothesis and the $z$-score are computed using eqs. (16) and (17), respectively. The quality of the fit is generally good for medium and low values of $g_{a\gamma}$, as can be seen from Figure 5, but it worsens for higher ones. We have checked in Appendix D that this does not significantly impact the main conclusions of this work.
Figure 5: Example distributions of the $\Pi$ statistic defined in eq. (19), comparing good and suboptimal performance of a classifier algorithm in different regions of the ALP parameter space
@@
Figure 6: ALP exclusion significance computed with the ML-based method, with the $2\sigma$ contours highlighted (solid lines). The contours obtained at the same confidence level with the LRT method are shown as dashed lines for comparison. A linear interpolation was applied between the points in the parameter space for which the exclusion confidence level was computed, which are indicated as black dots
@@ 5 Discussion of results
Our main results are presented in Figure 6, where a color map is used to represent the exclusion significance obtained with the ML classifier method in units of Gaussian standard deviations over the ALP parameter space. We note that, for both methods, the exclusion significance may be infinite for those points where the confidence level computed by eq. (16) is equal to 1. For plotting, these values are clipped to the highest finite significance; see also Appendix D.
@@
The $2\sigma$ sensitivity obtained with the method of ML classifiers, shown as solid lines in Figure 6, is comparable with that resulting from the LRT (dashed lines). For all sources, the overall shapes and extensions of the projected exclusion regions are in agreement between the two methods, with occasional low-significance “holes” or high-significance “islands” in both sets of contours. Such features are common to other analyses of this kind Ajello et al. (2016); Abdalla et al. (2021) and do not significantly affect our main results.
@@
The excluded region of the parameter space in the case of Mrk 501 is almost the same for both the baseline and flaring states. An explanation for this is the fact that the spectral shapes considered for the two states are fairly similar (see Fig. 1), and that the increased brightness of the flaring state is compensated by the longer exposure used for the baseline state. Given this, we can assume that the photon counts in the considered energy ranges are above the CTAO sensitivity for both states and are equally affected by ALPs, thus leading to similar exclusions in the parameter space. Conversely, PKS 2155-304 exhibits a spectral curvature and a stronger EBL absorption effect that significantly reduce the flux at higher energies. In this case, a flaring state can indeed bring the flux of higher-energy photons above the CTAO sensitivity, thus improving the constraining power of the observations. These qualitative considerations are confirmed by computing the Li-Ma significance $S$ of the signal Li and Ma (1983). While for Mrk 501 both observations yield $S\sim 300$, in the case of PKS 2155-304 a 5-hour flaring observation results in a significance $S\sim 800$, compared to $S\sim 200$ for a 50-hour observation in the baseline state.
@@
These projections, representing the mean expected CTAO sensitivity to ALPs from each considered source, are obtained by averaging the results from many different simulated observations. A different approach for estimating such a mean sensitivity, employed in various studies on ALPs Meyer et al. (2014); Meyer and Conrad (2014); Abdalla et al. (2021) and WIMPs Acharyya et al. (2021, 2023); Abe et al. (2024b), would be to use so-called “Asimov datasets” Cowan et al. (2011) instead. These datasets are generated by simply setting the numbers of counts in the ON and OFF regions equal to their expected values, in order to eliminate Poissonian fluctuations and directly provide a median estimate of the sensitivity. We have checked explicitly that the contours obtained from a single Asimov dataset using the LRT method are very close to those resulting from averaging over many “noisy” datasets. On the other hand, we found that applying the ML classifiers to Asimov datasets causes a significant overestimation of the exclusion regions, as the total absence of Poissonian noise in photon counts almost always produces low $p_{\rm ALP}$ values even for extremely small $g_{a\gamma}$. In a realistic analysis, these points of the parameter space would not be excluded, as the inevitable photon noise could easily mimic ALP effects at very low couplings. We thus conclude that Asimov datasets cannot be reliably used in estimating ALP sensitivities with ML classifiers.
@@
As a final remark, we note that these results may be affected by several systematic effects, such as the choice of energy binning, magnetic field parameters and EBL model. We discuss these effects in Appendix A. We note, in particular, that recent measurements of the cosmic infrared background by the CIBER experiment might extend the allowed parameter space for ALPs to a wider region than what is expected from the commonly adopted EBL models Kohri and Kodama (2017).
@@ 6 Conclusions
In this work, we have estimated the capability of CTAO to place upper limits on the ALP parameter space by simulating observations of the blazars Mrk 501 and PKS 2155-304 with the Northern and Southern Arrays of the Observatory. The source selection and observation criteria were motivated by the AGN Key Science Project described in ref. Acharya et al. (2018). Mock datasets for each source were prepared considering both long-term observations of baseline states, with spectra taken from the Fermi-LAT 4FGL-DR4 source catalog Ballet et al. (2024), and short flares based on previous TeV observations Albert et al. (2007); Aharonian et al. (2009).
@@
In Figure 7 we compare the projections derived in this work with previously published constraints on the ALP parameter space [^11] The reference limits are mostly taken from the online repository at ref. O’Hare (2020).. Most of these are computed at the $2\sigma$ confidence level Dessert et al. (2022); Noordhuis et al. (2023); Manzari et al. (2024); Ajello et al. (2016); Altenmüller et al. (2024); Abdalla et al. (2021); Abramowski et al. (2013), except for the limits by MAGIC which are reported at the $3\sigma$ level Abe et al. (2024a). In particular, ref. Abdalla et al. (2021) represents the only previous estimate of ALP sensitivity officially published by the CTAO Consortium, derived from simulated observations of the AGN NGC 1275 inside the Perseus galaxy cluster. With these targets, we expect that CTAO will be able to place limits on a much wider portion of previously unconstrained ALP parameter space in the $0.1\mbox{--}100\,\rm{neV}$ mass range. We finally remark that, while already competitive, the limits derived from a Southern source like PKS 2155-304 could be further improved with the foreseen addition of 2 LSTs to the Southern Array of CTAO, whose contribution would increase the sensitivity to low-energy spectral oscillations.
Figure 7: Comparison of the CTAO $2\sigma$ exclusion regions on the ALP parameter space obtained in this work with reference constraints from the literature (green). The projected sensitivities obtained from 50-hour observations of Mrk 501 (red) and PKS 2155-304 (blue) in their baseline states, as well as that from a 5-hour observation of a PKS 2155-304 flare (yellow) are shown. We do not explicitly include the results obtained from the Mrk 501 flare as they largely overlap with those from the baseline observation
Astrophysical ALP searches are affected by large systematic errors. These mainly stem from the uncertainty in the modeling of the morphology and intensity of the magnetic fields along the line of sight from the source. The magnetic field structure in blazar jets, in particular, is almost unconstrained experimentally and can only be estimated from theoretical arguments, thus making ALP searches from blazars somewhat model-dependent. In the absence of a more precise determination of this structure, turning to different computational approaches that could detect ALP-induced spectral features in a model-independent fashion might be a viable strategy. To probe this possibility, we computed CTAO sensitivity projections by applying the classical likelihood-ratio test and a newly proposed method making use of machine-learning classifiers. While in the present context the classifier-based method uses the same modeling assumptions as the traditional method, thus not really solving the issue of systematic uncertainties, a promising result is shown by the fact that the sensitivities obtained by the two results broadly agree with each other, as shown in Figure 6. This is a first step in the direction of exploring different non-parametric methods for ALP searches in astrophysical data. Future developments of this concept will involve dedicated studies focused on the application of more complex and optimized algorithms, such as neural networks, with the aim of maximizing the attainable experimental sensitivity to ALPs while reducing the amount of model-dependent assumptions.
@@ Acknowledgments
This work was conducted in the context of the CTAO DMEP working group. CTAO gratefully acknowledges financial support from the agencies and organizations listed at https://www.ctao.org/for-scientists/library/
acknowledgments/
. F. S. acknowledges financial support by the EU funding program “NextGenerationEU” in the context of the PNRR-IR “CTA+ - Cherenkov Telescope Array Plus” program.
@@
We would like to thank the computing centers that provided resources for the IRF generation, listed at
https://zenodo.org/records/5499840.
Appendix A Systematic effects
Figure 8a shows the average $2\sigma$ sensitivity on the ALP parameter space from a 50-hour observation of PKS 2155-304 in its baseline state for different choices of the energy binning. We compare the results presented in the main text, obtained for 20 bins per decade, with those resulting from considering either half or double the number of bins per decade (10 or 40, respectively). The latter choice of binning, in particular, was employed in ref. Abdalla et al. (2021). While there is very little difference between these results, halving the number of bins seems to reduce the size of the hole-like features around $g_{a\gamma}\sim 4\times 10^{-11}\,{\rm GeV}^{-1}$, while doubling it turns the feature into a continuous band. A similar outcome is found by comparing the results from different EBL models Franceschini et al. (2008); Finke et al. (2010); Gilmore et al. (2012) to the one used in the main text Domínguez et al. (2011), with the sensitivity to ALP parameters almost unaffected by this choice (Figure 8b).
Figure 8: CTAO sensitivity to the ALP parameter space obtained from a 50-hour simulated observation of PKS 2155-304 in its baseline state for different choices of the energy binning, EBL model, magnetic field strength and training strategy
Appendix B Details on spectral models
@@
The parameters of each spectral model considered in this work are reported in Table 3, where LP and ECPL refer to the log-parabola and power law with exponential cutoff models defined in eqs. (2) and (3). These models were implemented in Gammapy using the LogParabolaSpectralModel and ExpCutoffPowerLawSpectralModel classes, respectively. As discussed in the main text, the baseline spectra were taken from the Fermi-LAT 4FGL-DR4 catalog Ballet et al. (2024), while the flaring spectra were selected from previously published IACT observations. In particular, the EBL-corrected parameters of the Mrk 501 flaring spectrum are taken from Table 7 and Figure 19 of ref. Albert et al. (2007), referring to the flare of 30 June 2005. The PKS 2155-304 parameters are quoted from Table 3 in ref. Aharonian et al. (2009) and correspond to the data taken with the lowest energy threshold ($\rm 200\,GeV$, labelled as T200), corrected for EBL absorption.
Table 3: Parameters of the spectral models used for the sources considered in this work. We note that the $\beta$ parameter for the Mrk 501 flaring state is scaled by a factor $\ln(10)$ with respect to ref. Albert et al. (2007) after the conversion from base-10 logarithm to natural logarithm
Source Model $\phi_{0}$ $E_{0}$ $\alpha$ $\beta$ $E_{\rm cut}$ Ref. Baseline $10^{-12}\,\rm{MeV^{-1}\,cm^{-2}\,s^{-1}}$] [MeV] Mrk 501 LP $3.76\pm 0.05$ 1535.09 $1.75\pm 0.01$ $0.019\pm 0.004$ - Ballet et al. (2024) PKS 2155-304 LP $2.59\pm 0.13 11$60.97 $1.77\pm 0.01$ $0.042\pm 0.004$ - Ballet et al. (2024) Flaring $10^{-10}\,\rm{TeV^{-1}\,cm^{-2}\,s^{-1}}$] [TeV] [TeV] Mrk 501 LP $18.6\pm 0.06$ 0.3 $1.73\pm 0.06$ $0.13\pm 0.03$ - Albert et al. (2007) PKS 2155-304 ECPL $4.51\pm 0.36$ 1.0 $2.16\pm 0.05$ - $1.74\pm 0.24$ Aharonian et al. (2009)
@@
Figure 9: Accuracy of the XGBoost classifier grids defined over the ALP parameter space with the hyperparameters defined in Section 4, with the $90\%$ contours highlighted
Appendix C Classifier accuracy over the ALP parameter space